# Rethinking Aleatoric and Epistemic Uncertainty

### v2

@@ 5.3 Practical estimation of expected uncertainty reduction relies on approximations
Since we typically do not know the true data-generating process, $p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)$, a core approximation is to use a model over new data, $p_{n}(y^{\text{\scriptsize+}}_{1:m}|\pi)$, as a proxy. On top of this we might also need to approximate how our beliefs over $z$ update when we obtain new data. In principle the EUR is defined with respect to whichever updating scheme we are using, but in practice the true update can be too expensive to perform within an expectation over new data. This can be addressed by using some $q_{n+m}(z)$ in place of $p_{n+m}(z)$. A common approach is to assume $q_{n+m}(z)\propto p_{n}(y^{\text{\scriptsize+}}_{1:m}|z)p_{n}(z)$, even if the true update is not Bayesian (Bickford Smith et al,, 2023, 2024; Gal et al,, 2017; Kirsch et al,, 2019, 2023). It might be, for example, that we want to use a Bayesian model but cannot perform exact inference. If the true update is Bayesian and can be performed exactly then this assumption is of course not an approximation.
@@
$\mathrm{EUR}^{\prime}_{z}(\pi,m)=h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(y^{\text{\scriptsize+}}_{1:m}|\pi)}\!\left[h\!\left[q_{n+m}(z)\right]\right].$
The accuracy of this estimator depends on the mismatch between $p_{n}(y^{\text{\scriptsize+}}_{1:m}|\pi)$ and $p_{\mathrm{train}}(y^{\text{\scriptsize+}}_{1:m}|\pi)$ as well as the mismatch between $q_{n+m}(z)$ and $p_{n+m}(z)$, both of which are likely to be greater for larger $m$. Estimation therefore requires careful tradeoffs to mitigate these mismatches.
The practical relevance of this becomes clearer upon appreciating that the EUR estimator generalises a number of existing data-acquisition objectives. Under the exact-Bayesian-updating assumption it is equivalent to what has variously been called the “expected value of [additional/sample] information” (Bernardo & Smith,, 1994; Raiffa & Schlaifer,, 1961), the “expected $H_{\ell,\mathcal{A}}$-information gain” (Neiswanger et al,, 2022) and the “expected decision utility gain” (Huang et al,, 2024). From that objective we can then recover the expected information gain in $z$, which corresponds to the BALD score (Gal et al,, 2017; Houlsby et al,, 2011) if $z=\theta$ represents a set of stochastic model parameters, or the expected predictive information gain (Bickford Smith et al,, 2023) if $z=(x_{*},y_{*})$ represents a target input and its output. We note that $\mathrm{EUR}^{\prime}_{z}(\pi,m)$ is also closely related to the idea of the martingale posterior (Fong et al,, 2023) as $p_{n}(y^{\text{\scriptsize+}}_{1:m}|\pi)$ and our updating scheme together imply a joint distribution from which a martingale posterior can be derived.
Figure 3: Taking a decision-theoretic perspective allows us to disentangle model-based uncertainty from predictive performance and data dispersion. Here we have a model’s predictive distribution over an unknown variable, $z$, along with a reference distribution serving as a source of evaluation data. Using the model alone, we compute a measure of predictive uncertainty using a loss function, $\ell(a,z)$, where $a$ is an action. This differs from the data dispersion, which describes the reference distribution. Connecting the two distributions, the expected score (lower is better) measures the predictive performance of the model as judged using data from the reference distribution.
@@ 5.4 Model-based uncertainty should be used with care, and externally grounded evaluation is crucial
$=\mathbb{E}_{p_{\mathrm{eval}}(z)}\!\left[s(p_{n},z)\right]-h\!\left[p_{\mathrm{eval}}(z)\right]$ (5)
where $a^{*}_{\mathrm{eval}}$ is defined analogously to $a^{*}_{n}$ but with $p_{\mathrm{eval}}(z)$ as the predictive distribution. Equation 4 highlights that the discrepancy function measures the expected excess loss from acting based on $p_{n}(z)$ rather than $p_{\mathrm{eval}}(z)$ when $p_{\mathrm{eval}}(z)$ represents ground truth, which is connected to the idea of regret in learning theory (Szepesvári,, 2010). Equation 5 shows this is equivalent to the expected loss under $p_{\mathrm{eval}}(z)$ of the Bayes-optimal action derived from $p_{n}(z)$, minus the uncertainty measure from Section 5.1 applied to $p_{\mathrm{eval}}(z)$.
We can rearrange Equation 5 to highlight a decomposition:
$\underbrace{\mathbb{E}_{p_{\mathrm{eval}}(z)}\!\left[s(p_{n},z)\right]}_{\mathrm{expected\ score}}\,\,=\,\,\underbrace{d(p_{n},p_{\mathrm{eval}})}_{\mathrm{discrepancy}}\,\,+\underbrace{h\!\left[p_{\mathrm{eval}}(z)\right]}_{\mathrm{data\ dispersion}}$
where the expected score (lower is better) measures predictive performance and the data dispersion measures the spread in evaluation data drawn from the reference distribution. This generalises classic decompositions from statistics (Rice,, 2007) and information theory (Cover & Thomas,, 2005), as demonstrated in Examples 3 and 4.
Because past work has sometimes drawn connections between model-based uncertainty, predictive performance and data dispersion, we now explain how such connections can come about and emphasise that in the general case they should not be taken to hold. In particular, if we assume a specific model setup and estimation loss then we can derive Bayes estimators that generalise the predictive entropy and the expected conditional predictive entropy in Equation 1, but we emphasise that these assumptions will often not apply and that the estimators can be inaccurate.
@@
Three things are important to note in regard to Propositions 2 and 3. First, while they both suppose that $p_{n}(z)$ is designed to directly approximate $p_{\mathrm{eval}}(z)$, this need not always be the case. Because $p_{\mathrm{eval}}(z)$ might itself be imperfect, our evaluation might just be serving to provide a rough signal of the model’s predictive performance (Section 3.7). Second, they both assume a quadratic estimation loss, which might not reflect our preferences. An alternative loss would result in different Bayes estimators. For example, an absolute loss would lead us to use medians rather than expectations (Berger,, 1985). Third, they present estimators that are derived from subjective models and that therefore have no accuracy guarantee in the general case (Section 3.5).
Figure 4: BALD’s correspondence with infinite-step predictive information gain (a model’s long-run reduction in predictive entropy) can be weak, such that it is often better thought of as an estimator of a “true” one-step expected information gain in the model parameters. Here we show the behaviour of conjugate models trained on discrete (top) and continuous data (bottom). For $n\in(1,10,100,1000)$ we computed two estimation errors: $\varepsilon_{z}=(\mathrm{EIG}_{\theta}-\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty}))^{2}$ and $\varepsilon_{\theta}=(\mathrm{EIG}_{\theta}-\mathrm{EIG}_{\theta}^{\mathrm{true}})^{2}$. We show these errors (left; mean over 50 random-number seeds) along with the evolution of the predictive distribution (right). Both estimation errors are due to inaccurately simulating future data, an issue that resolves as $n$ increases and the predictive distribution converges to the data-generating process.
We therefore stress that model-based uncertainty should by default be considered separate from measures of predictive performance and data dispersion, with Figure 3 providing a concrete demonstration of their differences (see Appendix C for details). Uncertainty alone is not a reliable indicator of whether we can trust a model. Some kind of external grounding is crucial for well-informed practical deployment.
@@ 5.5 Popular information-theoretic quantities are best understood as imperfect estimators
$\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$ $=\mathrm{H}\!\left[p_{n}(z)\right]-\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right],$
measuring the reduction in the model’s predictive entropy from a Bayesian update on infinite new data, $y^{\text{\scriptsize+}}_{1:\infty}$.
@@
Let $y^{\text{\scriptsize+}}_{1:m}$ and $p_{n}(y|\theta)p_{n}(z|\theta)p_{n}(\theta)$ be a combination of data sequence and generative model that yield $p_{n}(\theta|y^{\text{\scriptsize+}}_{1:m})\to\delta_{\theta_{\infty}}(\theta)$ as $m\to\infty$. Then the expected conditional predictive entropy, $\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$, is a Bayes estimator of $\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right]$, the marginal predictive entropy after a Bayesian update on infinite new data, $y^{\text{\scriptsize+}}_{1:\infty}$.
@@
Assume the data and model from Proposition 4. Then the expected information gain in the model parameters, $\mathrm{EIG}_{\theta}$, from observing $z$ is a Bayes estimator of the infinite-step predictive information gain, $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$.
In Figure 4 we demonstrate that the approximation $\mathrm{EIG}_{\theta}\approx\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$ from Proposition 5 can be coarse. These results were produced with extremely simple setups within which we can perform exact inference and we are sure to recover the true data-generating process in the limit of infinite data (see Appendix C for details). We therefore know that the estimation error is due to a failure of the model to accurately simulate future data, which in turn is due to $n$ being finite.
This behaviour appears to align with existing results (with the caveat that past studies did not match the assumptions of Propositions 4 and 5). Figure 2 in Bickford Smith et al, (2024) and Figure 5 in Wimmer et al, (2023) show small-$n$ estimates of $\mathrm{EIG}_{\theta}$ that differ substantially from the changes in predictive entropy that actually occurred in practice. Those results even suggest it would have been more accurate to assume $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})=\mathrm{H}\!\left[p_{n}(z)\right]$, with $\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right]=0$, than to estimate it using $\mathrm{EIG}_{\theta}$. Meanwhile Mucsányi et al, (2024) and Valdenegro-Toro & Saromo-Mori, (2022) emphasised the “entanglement” of aleatoric- and epistemic-uncertainty estimators, which can be understood as the estimators themselves having an “epistemic” component—or, in our terminology, being inaccurate finite-data estimators.
@@
Returning to the same experimental setup as before, we find (Figure 4) that the approximation $\mathrm{EIG}_{\theta}\approx\mathrm{EIG}_{\theta}^{\mathrm{true}}$ is more accurate than $\mathrm{EIG}_{\theta}\approx\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$. In other words, BALD more closely tracks short-run changes in parameter uncertainty than it does long-run changes in predictive uncertainty. We do not claim this is a general result that will hold in all settings, but it is consistent with BALD being useful as a data-acquisition objective. The data-acquisition horizons in active learning are typically very short, so it is the short-run notion of information gain that matters, not the asymptotic notion. And while targeting predictions rather than parameters can be even more effective (Bickford Smith et al,, 2023, 2024), maximising short-run parameter information gain is still often preferable over random acquisition.
One takeaway here is that information-theoretic quantities should not be confused with the quantities they estimate. Another is that expected information gain in a variable of interest is a well motivated objective for data acquisition, assuming the action of interest is a probabilistic prediction of $z$ and we use a negative-log-likelihood loss function (Example 2), but it also crucially depends on the model’s ability to simulate future data, and it assumes a Bayesian update that might not match the true update (Section 5.3).
@@ 6 Conclusion
    BALD does not directly measure predictive-uncertainty reduction but is an (often inaccurate) estimator of it.
We suggest this new decision-theoretic view should be used in place of the predominant aleatoric-epistemic view as the field moves forward. Our hope is that this will support more productive discourse and methodological development.
A recurring point in our work is that true quantities of interest almost always have to be approximated in practice. Crucially a given practical quantity could be an accurate approximation of one true quantity and an inaccurate approximation of another. We stress that this dictates care in interpreting and using practical quantities. We also suggest that methods research should be guided by practical utility in concrete decision problems, such as data acquisition, rather than more abstract notions of approximation accuracy. Combined with stronger conceptual foundations, a renewed emphasis on real-world performance could accelerate progress in the use of uncertainty in machine learning.
@@ Acknowledgements
We thank Philip Dawid for sharing a technical report with us, and Andreas Kirsch and Mike Osborne for useful discussions and feedback. Freddie Bickford Smith is supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems (EP/L015897/1). Tom Rainforth is supported by EPSRC grant EP/Y037200/1.
@@ Appendix B Proofs of Propositions 4, 5 and 6
Let $y^{\text{\scriptsize+}}_{1:m}$ and $p_{n}(y|\theta)p_{n}(z|\theta)p_{n}(\theta)$ be a combination of data sequence and generative model that yield $p_{n}(\theta|y^{\text{\scriptsize+}}_{1:m})\to\delta_{\theta_{\infty}}(\theta)$ as $m\to\infty$. Then the expected conditional predictive entropy, $\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$, is a Bayes estimator of $\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right]$, the marginal predictive entropy after a Bayesian update on infinite new data, $y^{\text{\scriptsize+}}_{1:\infty}$.
@@
    Since $y^{\text{\scriptsize+}}_{1:\infty}$ recovers a single setting of $\theta$, reasoning about $\theta$ is equivalent to reasoning about $y^{\text{\scriptsize+}}_{1:\infty}$, following the argument presented in Fong et al, (2023). This allows us to apply Proposition 1 with $F=\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right]$ and $f(\theta)=\mathrm{H}\!\left[p_{n}(z|\theta)\right]$, which gives $\eta^{*}=\mathbb{E}_{p_{n}(\theta)}\!\left[\mathrm{H}\!\left[p_{n}(z|\theta)\right]\right]$. ∎
@@
Assume the data and model from Proposition 4. Then the expected information gain in the model parameters, $\mathrm{EIG}_{\theta}$, from observing $z$ is a Bayes estimator of the infinite-step predictive information gain, $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$.
@@
    The information gain to be estimated, $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$, is defined as the reduction in the model’s predictive entropy from a Bayesian update on infinite new data, $y^{\text{\scriptsize+}}_{1:\infty}$:
    $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$ $=\mathrm{H}\!\left[p_{n}(z)\right]-\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right].$
    Combining the known $\mathrm{H}\!\left[p_{n}(z)\right]$ with the Bayes estimator of $\mathrm{H}\!\left[p_{n}(z|y^{\text{\scriptsize+}}_{1:\infty})\right]$ from Proposition 4 gives
@@
    as a Bayes estimator of $\mathrm{IG}_{z}(y^{\text{\scriptsize+}}_{1:\infty})$. ∎

### v3

@@ 5.3 Practical estimation of expected uncertainty reduction relies on approximations
Since we typically do not know the true data-generating process, $p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{1:m}|\pi)$, a core approximation is to use a model over new data, $p_{n}(y^{\scriptscriptstyle+}_{1:m}|\pi^{\prime})$, as a proxy. On top of this we might also need to approximate how our beliefs over $z$ update when we obtain new data. In principle the EUR is defined with respect to whichever updating scheme we are using, but in practice the true update can be too expensive to perform within an expectation over new data. This can be addressed by using some $q_{n+m}(z)$ in place of $p_{n+m}(z)$. A common approach is to assume a Bayesian belief update given $y^{\scriptscriptstyle+}_{1:m}$, even if the true update is not Bayesian (Bickford Smith et al,, 2023, 2024; Gal et al,, 2017; Kirsch et al,, 2019, 2023). It might be, for example, that we want to use a Bayesian model but cannot perform exact inference. If the true update is Bayesian and can be performed exactly then this assumption is of course not an approximation.
@@
$\mathrm{EUR}^{\mathrm{est}}_{z}(\pi^{\prime},m)=h\!\left[p_{n}(z)\right]-\mathbb{E}_{p_{n}(y^{\scriptscriptstyle+}_{1:m}|\pi^{\prime})}\!\left[h\!\left[q_{n+m}(z)\right]\right].$
The accuracy of this estimator depends on the mismatch between $p_{n}(y^{\scriptscriptstyle+}_{1:m}|\pi^{\prime})$ and $p_{\mathrm{train}}(y^{\scriptscriptstyle+}_{1:m}|\pi)$ as well as the mismatch between $q_{n+m}(z)$ and $p_{n+m}(z)$, both of which are likely to be greater for larger $m$. Estimation therefore requires careful tradeoffs to mitigate these mismatches.
The practical relevance of this becomes clearer upon appreciating that the EUR estimator generalises a number of existing data-acquisition objectives. Under the exact-Bayesian-updating assumption it is equivalent to what has variously been called the “expected value of [additional/sample] information” (Bernardo & Smith,, 1994; Raiffa & Schlaifer,, 1961), the “expected $H_{\ell,\mathcal{A}}$-information gain” (Neiswanger et al,, 2022) and the “expected decision utility gain” (Huang et al,, 2024). From that objective we can then recover the expected information gain in $z$, which corresponds to the BALD score (Gal et al,, 2017; Houlsby et al,, 2011) if $z=\theta$ represents a set of stochastic model parameters, or the expected predictive information gain (Bickford Smith et al,, 2023) if $z=(x_{*},y_{*})$ represents a target input and its output. Notably $\mathrm{EUR}^{\mathrm{est}}_{z}(\pi^{\prime},m)$ is also closely related to the idea of the martingale posterior (Fong et al,, 2023) as $p_{n}(y^{\scriptscriptstyle+}_{1:m}|\pi^{\prime})$ and our updating scheme together imply a joint distribution from which a martingale posterior can be derived.
Figure 3: Model-based uncertainty, predictive performance and data dispersion are related but distinct quantities. Here we show a model’s predictive distribution over an output, $z\in\mathbb{R}$, corresponding to an input, $x\in\mathbb{R}$; we also show a reference distribution serving as a source of evaluation data. Using the model alone, we compute a measure of predictive uncertainty using a loss function, $\ell(a,z)$, where $a$ is an action. This differs from the data dispersion, which describes the reference distribution. Connecting the two distributions, the expected score (lower is better) measures the predictive performance of the model as judged using data from the reference distribution.
@@ 5.4 Model-based uncertainty should be used with care, and externally grounded evaluation is crucial
$=\underbrace{\mathbb{E}_{p_{\mathrm{eval}}(z)}\!\left[s(p_{n},z)\right]}_{\mathrm{expected\ score}}\,\,-\underbrace{h\!\left[p_{\mathrm{eval}}(z)\right]}_{\mathrm{data\ dispersion}}$ (5)
where $a^{*}_{\mathrm{eval}}$ is defined analogously to $a^{*}_{n}$ but with $p_{\mathrm{eval}}(z)$ as the predictive distribution. Equation 4 highlights that the discrepancy is the expected excess loss from acting based on $p_{n}(z)$ rather than $p_{\mathrm{eval}}(z)$ when $p_{\mathrm{eval}}(z)$ represents the ground truth, which is linked to the idea of regret (Szepesvári,, 2010). Equation 5 shows that the discrepancy also corresponds to the expected score of $p_{n}(z)$, which measures predictive performance (lower is better), minus the uncertainty measure from Section 5.1 applied to $p_{\mathrm{eval}}(z)$, which measures the statistical dispersion in evaluation data drawn from the reference distribution. This relationship generalises classic decompositions from statistics (Rice,, 2007) and information theory (Cover & Thomas,, 2005), as shown in Examples 3 and 4. Notably the expected score can be seen as a form of frequentist risk for an “outer” decision problem; another form also averages over the training data.
Because past work has sometimes drawn connections between model-based uncertainty, predictive performance and data dispersion, we now explain how such connections can come about and why in the general case they should not be taken to hold. In particular, if we assume a specific model setup and estimation loss then we can derive Bayes estimators that generalise the marginal predictive entropy and the expected conditional predictive entropy in Equation 1, but these assumptions will typically not apply and, even if they do apply, the estimators can be highly inaccurate.
@@
Three things are important to note in regard to Propositions 2 and 3. First, while they both suppose that $p_{n}(z)$ is designed to directly approximate $p_{\mathrm{eval}}(z)$, this need not always be the case. Because $p_{\mathrm{eval}}(z)$ might itself be imperfect, our evaluation might just be serving to provide a rough signal of the model’s predictive performance (Section 3.7). Second, they both assume a quadratic estimation loss, which might not reflect our preferences. An alternative loss would result in different Bayes estimators. For example, an absolute loss would lead us to use medians rather than expectations (Berger,, 1985). Third, they present estimators that are derived from subjective models and that, as a result, have no accuracy guarantee in the general case (Section 3.5).
Figure 4: BALD’s correspondence with infinite-step predictive information gain (a model’s long-run reduction in predictive entropy) can be weak, such that it is often better thought of as an estimator of a “true” one-step expected information gain in the model parameters. Here we show the behaviour of conjugate models trained on discrete (top) and continuous data (bottom). For $n\in(1,10,100,1000)$ we computed two estimation errors: $\varepsilon_{z}=(\mathrm{EIG}_{\theta}-\mathrm{IG}_{z}(y^{\scriptscriptstyle+}_{1:\infty}))^{2}$ and $\varepsilon_{\theta}=(\mathrm{EIG}_{\theta}-\mathrm{EIG}_{\theta}^{\mathrm{true}})^{2}$. We show these errors (left; mean over 50 random-number seeds) along with the evolution of the predictive distribution (right). Both estimation errors are due to inaccurately simulating future data, an issue that resolves as $n$ increases and the predictive distribution converges to the data-generating process.
We therefore stress that model-based uncertainty should be considered separate from predictive performance and data dispersion, as shown in Figure 3 (see Appendix C for details). Uncertainty alone is not a reliable indicator of whether we can trust a model. Some kind of external grounding is crucial for well-informed practical deployment.
@@ 5.5 Popular information-theoretic quantities are best understood as imperfect estimators
[…]
measuring the reduction in the model’s predictive entropy from a Bayesian update on infinite new data, $y^{\scriptscriptstyle+}_{1:\infty}$.
@@
[…]
@@
[…]
This behaviour appears to align with existing results (with the caveat that past studies did not match the assumptions of Propositions 4 and 5). Figure 2 in Bickford Smith et al, (2024) and Figure 5 in Wimmer et al, (2023) show small-$n$ estimates of $\mathrm{EIG}_{\theta}$ that differ substantially from the changes in predictive entropy that actually occurred in practice. Those results even suggest it would have been more accurate to assume $\mathrm{IG}_{z}(y^{\scriptscriptstyle+}_{1:\infty})=\mathrm{H}\!\left[p_{n}(z)\right]$, with $\mathrm{H}\!\left[p_{n}(z|y^{\scriptscriptstyle+}_{1:\infty})\right]=0$, than to estimate it using $\mathrm{EIG}_{\theta}$. Meanwhile Mucsányi et al, (2024) and Valdenegro-Toro & Saromo-Mori, (2022) emphasised the “entanglement” of aleatoric- and epistemic-uncertainty estimators, which can be understood as the estimators themselves having an “epistemic” component—or, in our terminology, being inaccurate finite-data estimators.
@@
[…]
One takeaway here is that information-theoretic quantities should not be confused with the quantities they estimate. Another is that expected information gain in a variable of interest is a well-motivated objective for data acquisition, assuming the action of interest is a probabilistic prediction of $z$ and we use a negative-log-likelihood loss function (Example 2), but it also crucially depends on the model’s ability to simulate future data, and it assumes a Bayesian update that might not match the true update (Section 5.3).
@@ 6 Conclusion
    BALD does not directly measure long-run reducible predictive uncertainty but rather estimates it, and the associated estimation error can be large.
We believe our decision-theoretic view should be used in place of the predominant aleatoric-epistemic view as the field moves forward. Our hope is that it will support more productive discourse and methodological development.
We also urge care in using uncertainty and related quantities in practice. A crucial recurring point in this work is that true quantities of interest almost always have to be approximated in the real world. Ignoring approximation errors can lead to dangerous assumptions and ineffective methods.
Finally we suggest that future methods research should be guided by practical utility in concrete decision problems, such as data acquisition and model selection, rather than more abstract notions of approximation accuracy. Combined with stronger conceptual foundations, a renewed emphasis on real-world performance could accelerate progress in the use of probabilistic reasoning in machine learning.
@@ Acknowledgements
We thank Philip Dawid for sharing a technical report with us. We are also grateful to Andreas Kirsch, Mike Osborne and the anonymous reviewers of this paper for useful discussions and feedback. Freddie Bickford Smith is supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems (EP/L015897/1). Tom Rainforth is supported by EPSRC grant EP/Y037200/1.
@@ Appendix B Proofs of Propositions 4, 5 and 6
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]